# Your Student is Better Than Expected: Adaptive Teacher-Student Collaboration for Text-Conditional Diffusion Models

### 17 Dec 2023 v1

@@ Abstract
Knowledge distillation methods have recently shown to be a promising direction to speedup the synthesis of large-scale diffusion models by requiring only a few inference steps. While several powerful distillation methods were recently proposed, the overall quality of student samples is typically lower compared to the teacher ones, which hinders their practical usage. In this work, we investigate the relative quality of samples produced by the teacher text-to-image diffusion model and its distilled student version. As our main empirical finding, we discover that a noticeable portion of student samples exhibit superior fidelity compared to the teacher ones, despite the “approximate” nature of the student. Based on this finding, we propose an adaptive collaboration between student and teacher diffusion models for effective text-to-image synthesis. Specifically, the distilled model produces the initial sample, and then an oracle decides whether it needs further improvements with a slow teacher model. Extensive experiments demonstrate that the designed pipeline surpasses state-of-the-art text-to-image alternatives for various inference budgets in terms of human preference. Furthermore, the proposed approach can be naturally used in popular applications such as text-guided image editing and controllable generation.
@@ 1 Introduction
There are two major research directions mitigating the sequential inference problem of state-of-the-art diffusion models. One of them considers the inference process as a solution of a probability flow ODE and designs efficient and accurate solvers [49, 18, 26, 27, 58] reducing the number of inference steps down to ${\sim}10$ without drastic loss in image quality. Another direction represents a family of knowledge distillation approaches [42, 50, 30, 24, 25, 28, 12, 44] that learn the student model to simulate the teacher distribution requiring only $1{-}4$ inference steps. Recently, distilled text-to-image models have made a significant step forward [30, 25, 28, 44]. However, they still struggle to achieve the teacher performance either in terms of image fidelity and textual alignment [30, 25, 28] or distribution diversity [44]. Nevertheless, we hypothesize that text-to-image students may already have qualitative advantages over their teachers. If so, perhaps it would be more beneficial to consider a student-teacher collaboration rather than focusing on replacing the teacher model entirely.
In this paper, we take a closer look at images produced by distilled text-conditional diffusion models and observe that the student can generate some samples even better than the teacher. Surprisingly, the number of such samples is significant and sometimes reaches up to half of the empirical distribution. Based on this observation, we design an adaptive collaborative pipeline that leverages the superiority of student samples and outperforms both individual models alone for various inference budgets. Specifically, the student model first produces the initial image sample given the text prompt, then the “oracle” decides if this sample should be further refined with the teacher model. The similar idea has recently demonstrated its effectiveness for large language models (LLMs) [5] and we show that it can be naturally extended for text-conditional diffusion models as well. Our approach is schematically presented in Figure 1. To summarize, our paper presents the following contributions:
@@
  • Based on the finding above, we develop an adaptive teacher-student collaborative approach for cost-effective text-to-image synthesis. The method not only increases the inference efficiency but also improves the generative quality by exploiting the superior student samples.
@@ 2 Related work
Along with cascaded models, there are other works combining several diffusion models into a single pipeline. Some methods propose using distinct diffusion models at different time steps [1, 23, 8, 55]. Others  [34, 25] consider an additional model to refine the samples produced with a main model. In contrast, our method relies on the connection between student and teacher models and adaptively improves only selected student samples to reduce the inference costs.
@@ 3 Toward a unified teacher-student framework
Opposed to the purpose of replacing the expensive text-to-image diffusion models by more effective few-step alternatives, the present work suggests considering the distilled text-to-image models as a firm companion in teacher-student collaboration.
@@ 3.1 Delving deeper into the student performance
Figure 3: (a) Visual examples of similar (Left) and dissimilar (Right) teacher and student samples. (b) Similarity between the student and teacher samples w.r.t. the difference in sample quality. Highly distinct samples tend to be of different quality. (c) Human vote distributions for different distance ranges between student and teacher samples. Most student wins are achieved when the student diverges from the teacher.
Figure 4: Effect of image complexity. (a) More similar student and teacher samples corresponds to simpler images and vice versa. (b) The student and teacher largely diverge in image quality on the complex teacher samples. Notably, the student often outperforms the teacher on these samples.
@@
Primarily, we observe that many student samples are highly distinct from the teacher ones. A few image pairs are presented in Figure 3a. Figure 3c presents the human vote distribution over three distance ranges: low, medium, and high. Interestingly, most student wins appear when its samples are highly different from the teacher ones. This brings us to our second observation:
Also, we evaluate the relative gap in sample quality against the similarity between the teacher and student outputs. To measure the quality of individual samples, we use ImageReward [53], which shows a positive correlation with human preferences in terms of image fidelity and prompt alignment. The gap in sample quality is calculated as the difference between the ImageReward scores for student and teacher samples. We observe that highly distinct samples likely have a significant difference in quality. Importantly, this holds for both student failures and successes, as shown in Figure 3b. Therefore, by effectively detecting the positive student samples and improving the negative ones, we can potentially increase the overall generative performance.
@@
Text prompts. Then, we analyse the connection of the student-teacher similarity with the prompt length. Figure 5 demonstrates that shorter prompts typically lead to more similar teacher and student samples. Here, the prompt length equals to the number of CLIP tokens. Intuitively, longer prompts are more likely to describe intricate scenes and object compositions than shorter ones. Note that long prompts may carry various textual informativeness and describe concepts of different complexity. We hypothesize that this causes high variance in student-teacher similarity.
Also, we report the prompt influence on the student generation w.r.t. the student-teacher similarity in Figure 5b. We estimate the prompt influence by aggregating student cross-attention maps. More details are in B. The student tends to imitate the teacher if it relies heavily on the text prompt.
Trajectory curvature. Previously, it was shown to be beneficial to straighten the PF-ODE trajectory before distillation [24, 25]. We investigate the effect of the trajectory curvature on the similarity between the teacher and student samples and its correlation with the teacher sample complexity. We calculate the trajectory curvatures following  [4] and observe that straighter trajectories lead to more similar student and teacher samples (Figure 6,a). In addition, we show, for the first time, that the trajectory curvature correlates with the teacher sample complexity (Figure 6,b).
@@
To sum up, we conclude that the student largely diverges from the teacher on the samples that are challenging in different respects. Interestingly, the student can regularly outperform the teacher in these cases.
@@ 3.2 Method
Figure 9: User preference study (SD1.5). (a) Comparison of our approach to the top-performing teacher configurations. (b) Comparison to the teacher model with DPM-Solver for the same number of steps. (c) Comparison to the refinement strategy without the adaptive step for the same number of steps. Top row: LAION-Aesthetic text prompts. Bottom row: COCO2014 text prompts. For our adaptive approach, we use the refinement strategy (R).
@@ 4.1 Text-guided image synthesis
Also, we evaluate user preferences using side-by-side comparisons conducted by professional assessors. We select random $600$ text prompts from the COCO2014 validation set and $600$ prompts from LAION-Aesthetics [^2] https://laion.ai/blog/laion-aesthetics/. More details on the human evaluation pipeline are provided in D.1.
Configuration. For our adaptive approach, we consider both refinement (R) and regeneration (G) strategies using a second order multistep DPM solver [27] and vary the number of sampling steps depending on the inference budget. As a sample estimator $\mathbf{E}$, we consider ImageReward, except for the CLIP score evaluation. For each inference budget, we tune the hyperparameters $\sigma$ and $\tau$ on the hold-out prompt set. The exact values are provided in D.2.
@@
Figure 11: User preference study (SDXL). Left: Comparison of the adaptive approach with CD-SDXL to the top-performing teacher setup. Right: Comparison of the adaptive approach with ADD-XL to ADD-XL for the same number of sampling steps.
Figure 12: User preferences for the different accuracy levels of the sample estimator. Current state represents the results with the ImageReward estimator. Other results demonstrate the potential gains if the estimator performance improves in the future.
@@
Effect of estimator accuracy. The potential bottleneck of our approach is the poor correlation of existing text-to-image automated estimators with human preferences. For example, ImageReward usually exhibits up to $65\%$ correlation with the annotators. Here, we investigate the effect of the estimator accuracy on the performance of our scheme to reveal its potential if more accurate estimators are developed in future work. Figure 12 presents our method performance for different estimator accuracy rates and demonstrates significant future gains even for the $75\%$ correlation.
SDXL results. In addition to SD1.5 experiments, we evaluate our pipeline using the recent CD-SDXL [28] and ADD-XL [44] which are both distilled from the SDXL-Base model [34]. Our approach with CD-SDXL stands against the top-performing SDXL setting: $50$ steps of the DDIM sampler. For ADD-XL, we provide the comparison for $4$ steps where ADD-XL has demonstrated exceptionally strong generative performance in terms of human preference [44]. In both settings, our approach uses the adaptive refinement strategy with UniPC solver [58]. Note that both the SDXL-Base and SDXL-Refiner [34] models can be used for refinement. In our experiments, we observe that the refiner suits slightly better for fixing minor defects while the teacher allows more pronounced changes. Thus, we use the refiner for low $\sigma$ values and the base model for the higher ones. More setup details are provided in D.3.
@@ 5 Conclusion
This work investigates the performance of the distilled text-to-image models and demonstrates that they may consistently outperform the teachers on many samples. We design an adaptive text-to-image generation pipeline that takes advantage of successful student samples and, in combination with the teacher model, outperforms other alternatives for low and high inference budgets. Importantly, text-to-image automatic estimators and distillation methods still have room for further development. Therefore, we believe the proposed approach may be even in higher demand when future works make a significant step forward in these directions.
@@ Appendix C Cut-off threshold tuning
The cut-off threshold $\tau$ is used for the adaptive selection of student samples. It corresponds to the $k$-th percentile of the metric values calculated on validation student samples. $600$ and $300$ samples are generated for tuning on the COCO2014 and LAION-Aesthetics datasets, respectively. Note that the prompts used for tuning do not overlap with the test ones. Then, we calculate an individual score, e.g., IR score, for each validation student sample and select the percentile based on an inference budget or target metric value. For example, suppose we select the percentile given a $15$ step budget and intend to perform $5$ student steps and $20$ steps for the improvement strategy. In this case, we have to select $\tau$ as a $50$-th percentile, which results in the final number of steps: $5+0.5\cdot 20=15$.

### 28 Dec 2023 v2

@@ Abstract
Knowledge distillation methods have recently shown to be a promising direction to speedup the synthesis of large-scale diffusion models by requiring only a few inference steps. While several powerful distillation methods were recently proposed, the overall quality of student samples is typically lower compared to the teacher ones, which hinders their practical usage. In this work, we investigate the relative quality of samples produced by the teacher text-to-image diffusion model and its distilled student version. As our main empirical finding, we discover that a noticeable portion of student samples exhibit superior fidelity compared to the teacher ones, despite the “approximate” nature of the student. Based on this finding, we propose an adaptive collaboration between student and teacher diffusion models for effective text-to-image synthesis. Specifically, the distilled model produces an initial image sample, and then an oracle decides whether it needs further improvements with the teacher model. Extensive experiments demonstrate that the designed pipeline surpasses state-of-the-art text-to-image alternatives for various inference budgets in terms of human preference. Furthermore, the proposed approach can be naturally used in popular applications such as text-guided image editing and controllable generation.
@@ 1 Introduction
There are two major research directions mitigating the sequential inference problem of state-of-the-art diffusion models. One of them considers the inference process as a solution of a probability flow ODE and designs efficient and accurate solvers [49, 18, 26, 27, 58] reducing the number of inference steps down to ${\sim}10$ without drastic loss in image quality. Another direction represents a family of knowledge distillation approaches [42, 50, 30, 24, 25, 28, 12, 44] that learn the student model to simulate the teacher distribution requiring only $1{-}4$ inference steps. Recently, distilled text-to-image models have made a significant step forward [30, 25, 28, 44]. However, they still struggle to achieve the teacher performance either in terms of image fidelity and textual alignment [30, 25, 28] or distribution diversity [44]. Nevertheless, we hypothesize that text-to-image students may already have qualitative merits over their teachers. If so, perhaps it would be more beneficial to consider a teacher-student collaboration rather than focusing on replacing the teacher model entirely.
In this paper, we take a closer look at images produced by distilled text-conditional diffusion models and observe that the student can generate some samples even better than the teacher. Surprisingly, the number of such samples is significant and sometimes reaches up to half of the empirical distribution. Based on this observation, we design an adaptive collaborative pipeline that leverages the superiority of student samples and outperforms both individual models alone for various inference budgets. Specifically, the student model first generates an initial image sample given a text prompt, and then an “oracle” decides if this sample should be updated using the teacher model at extra compute. The similar idea has recently demonstrated its effectiveness for large language models (LLMs) [5] and we show that it can be naturally applied to text-conditional diffusion models as well. Our approach is schematically presented in Figure 1. To summarize, our paper presents the following contributions:
@@
  • Based on the finding above, we develop an adaptive teacher-student collaborative approach for effective text-to-image synthesis. The method not only reduces the average inference costs but also improves the generative quality by exploiting the superior student samples.
@@ 2 Related work
Along with cascaded models, there are other works combining several diffusion models into a single pipeline. Some methods propose to use distinct diffusion models at different time steps [1, 23, 8, 55]. Others [34, 25] consider an additional model to refine the samples produced with a base model. In contrast, our method relies on the connection between student and teacher models and adaptively improves only selected student samples to reduce the inference costs.
@@ 3 Toward a unified teacher-student framework
Opposed to the purpose of replacing the expensive text-to-image diffusion models by more effective few-step alternatives, the present work suggests considering the distilled text-to-image models as a firm companion in a teacher-student collaboration.
@@ 3.1 Delving deeper into the student performance
Figure 3: (a) Visual examples of similar (Left) and dissimilar (Right) teacher and student samples. (b) Similarity between the student and teacher samples w.r.t. the difference in sample quality. Highly distinct samples tend to be of different quality. (c) Human vote distribution for different distance ranges between student and teacher samples. Most of the student wins are achieved when the student diverges from the teacher.
Figure 4: Effect of image complexity. (a) More similar student and teacher samples corresponds to simpler images and vice versa. (b) The student and teacher largely diverge in image quality on the complex teacher samples.
@@
Primarily, we observe that many student samples are highly distinct from the teacher ones. A few image pairs are presented in Figure 3a. Figure 3c presents the human vote distribution for three distance ranges: low, medium, and high. Interestingly, most of the student wins appear when its samples are highly different from the teacher ones. This brings us to our second observation:
The student wins are more likely where its samples significantly differ from the teacher ones.
Also, we evaluate the relative gap in sample quality against the similarity between the teacher and student outputs. To measure the quality of individual samples, we use ImageReward [53], which shows a positive correlation with human preferences in terms of image fidelity and prompt alignment. The gap in sample quality is calculated as the difference between the ImageReward scores for student and teacher samples. We observe that highly distinct samples likely have a significant difference in quality. Importantly, this holds for both student failures and successes, as shown in Figure 3b. Therefore, effectively detecting the positive student samples and improving the negative ones can potentially increase the generative performance.
@@
Text prompts. Then, we analyse the connection of the student-teacher similarity with the prompt length. Figure 5 demonstrates that shorter prompts typically lead to more similar teacher and student samples. Here, the prompt length equals to the number of CLIP tokens. Intuitively, longer prompts are more likely to describe intricate scenes and object compositions than shorter ones. Note that long prompts can also carry low textual informativeness and describe concepts of low complexity. We hypothesize that this causes high variance in Figure 5a.
Also, we report the prompt influence on the student generation w.r.t. the student-teacher similarity in Figure 5b. We estimate the prompt influence by aggregating student cross-attention maps. More details are in B.1. The student tends to imitate the teacher if it relies heavily on the text prompt.
Trajectory curvature. Previously, it was shown to be beneficial to straighten the PF-ODE trajectory before distillation [24, 25]. We investigate the effect of the trajectory curvature on the similarity between the teacher and student samples and its correlation with the teacher sample complexity. We estimate the trajectory curvatures following [4] and observe that straighter trajectories lead to more similar student and teacher samples (Figure 6a). In addition, we show that the trajectory curvature correlates with the teacher sample complexity (Figure 6b).
@@
To sum up, we conclude that the student largely diverges from the teacher on the samples that are challenging in different respects. Interestingly, the superior student samples often occur in these cases.
@@ 3.2 Method
Figure 9: User preference study (SD1.5). (a) Comparison of our approach to the top-performing teacher configurations. (b) Comparison to the teacher model with DPM-Solver for the same average number of steps. (c) Comparison to the refinement strategy without the adaptive step for the same average number of steps. Top row: LAION-Aesthetic text prompts. Bottom row: COCO2014 text prompts. For our adaptive approach, we use the refinement strategy (R).
@@ 4.1 Text-guided image synthesis
Also, we evaluate user preferences using side-by-side comparisons conducted by professional assessors. We select $600$ random text prompts from the COCO2014 validation set and $600$ from LAION-Aesthetics [^2] https://laion.ai/blog/laion-aesthetics/. More details on the human evaluation pipeline are provided in D.1.
Configuration. For our adaptive approach, we consider both refinement (R) and regeneration (G) strategies using a second order multistep DPM solver [27] and vary the number of sampling steps depending on the average inference budget. As a sample estimator $\mathbf{E}$, we consider ImageReward, except for the CLIP score evaluation. For each inference budget, we tune the hyperparameters $\sigma$ and $\tau$ on the hold-out prompt set. The exact values are provided in D.2.
@@
Figure 11: User preference study (SDXL). Left: Comparison of the adaptive approach with CD-SDXL to the top-performing teacher setup. Right: Comparison of the adaptive approach with ADD-XL to ADD-XL for the same average number of steps.
Figure 12: User preferences for different accuracy levels of the no-reference decision-making procedure. Current state represents the results using ImageReward. The results for higher accuracy rates demonstrate the future gains if the oracle performance improves.
@@
Effect of oracle accuracy. The potential bottleneck of our approach is a poor correlation of existing text-to-image automated estimators with human preferences. For example, ImageReward usually exhibits up to $65\%$ agreement with annotators. Moreover, it remains unclear what oracle accuracy can be achieved with no-reference decision-making, even if the estimator provides the perfect agreement.
In Figure 12, we conduct a synthetic experiment examining the effect of the oracle accuracy on our scheme performance to reveal its future potential. We compare the adaptive refinement method ($10$ steps) to SD1.5 ($50$ steps) manually varying the oracle accuracy. We observe significant future gains even for the $75\%$ accuracy rate.
SDXL results. In addition to SD1.5 experiments, we evaluate our pipeline using the recent CD-SDXL [28] and ADD-XL [44] which are both distilled from the SDXL-Base model [34]. Our approach with CD-SDXL stands against the top-performing SDXL setting: $50$ steps of the DDIM sampler. For ADD-XL, we provide the comparison for $4$ steps where ADD-XL has demonstrated exceptionally strong generative performance in terms of human preference [44]. In both settings, our approach uses the adaptive refinement strategy with the UniPC solver [58]. Note that both the SDXL-Base and SDXL-Refiner [34] models can be used for refinement. In our experiments, we observe that the refiner suits slightly better for fixing minor defects while the teacher allows more pronounced changes. Thus, we use the refiner for low $\sigma$ values and the base model for the higher ones. More setup details are provided in D.3.
@@ 5 Conclusion
This work investigates the performance of the distilled text-to-image models and demonstrates that they may consistently outperform the teachers on many samples. We design an adaptive text-to-image generation pipeline that takes advantage of successful student samples and, in combination with the teacher model, outperforms other alternatives for low and high inference budgets. Importantly, text-to-image automated estimators and distillation methods still have room for further development. Therefore, we believe the proposed approach may be even in higher demand when future works make a significant step forward in these directions.
@@ Appendix C Cut-off threshold tuning
The cut-off threshold $\tau$ is used for the adaptive selection of student samples. It corresponds to the $k$-th percentile of the metric values calculated on validation student samples. $600$ and $300$ samples are generated for tuning on the COCO2014 and LAION-Aesthetics datasets, respectively. Note that the prompts used for tuning do not overlap with the test ones. Then, we calculate an individual score, e.g., IR score, for each validation student sample and select the percentile based on an average inference budget or target metric value. For example, suppose we select the percentile given a $15$ step budget and intend to perform $5$ student steps and $20$ steps for the improvement strategy. In this case, we have to select $\tau$ as a $50$-th percentile, which results in the final average number of steps: $5+0.5\cdot 20=15$.